A 3.11 mol sample of CH4 (Methane) gas has a volume of 329 mL at a temperature of 79°C.
i. Calculate the pressure using the Ideal Gas Law.
ii. Calculate the corrected pressure using the Van der Waals Equation.
atm×L2 L a=2.268 mol2 b=0.04301 mol
1
Expert's answer
2020-04-19T07:08:22-0400

i. p*V = n*R*T

p is in atm, V is in L, n is in moles, R = 0.082 L*atm/(K*mole), T in K.

p = 3.11*0.082*(79+273)/0.329 = 272.85 atm


ii. (p + a*n2/V2)*(V - n*b) = n*R*T

p = 3.11*0.082*(79+273)/(0.329 - 3.11*0.04301) - 2.268*(3.11)2/(0.329)2 = 257.12 atm